Organizing Committee of The Symposium on Ring Theory and Representation Theory

The symposium on Ring Theory and Representation Theory has been held annually in Japan and the Proceedings have been published by the organizing committee. The first Symposium was organized in 1968 by H. Tominaga, H. Tachikawa, M. Harada and S. Endo. After their retirement, a new committee was organized in 1997 for managing the Symposium. The present members of the committee are Y. Hirano (Okayama Univ.), S. Koshitani (Chiba Univ.), K. Nishida (Shinshu Univ.) and M. Sato (Yamanashi Univ.).

The Proceedings of each Symposium is edited by program organizer. Anyone who wants these Proceedings should ask to the program organizer of each Symposium or one of the committee members.

The Symposium in 2006 will be held at Hiroshima University in Hiroshima Prefecture for Sep 16-18, and the program will be arranged by Mamoru Kutami (Ymaguchi Univ.).

Concerning several information on ring theory group in Japan containing schedules of meetings and symposiums as well as the addresses of members in the group, you should refer the following homepage, which is arranged by M. Sato (Yamanashi Univ.):

http://fuji.cec.yamanashi.ac.jp/ ring/ (in Japanese) civil2.cec.yamanashi.ac.jp/ ring/japan/ (in English)

Kenji Nishida Matsumoto, Japan December, 2005